OpenStudy (anonymous):

A few algebra questions (sigmas, sequences, etc.) (Will post questions in the comments)

OpenStudy (anonymous):

"Which represents the series using sigma notation?"\[1, \frac{ 1 }{ 3 }, \frac{ 1 }{ 9 }, \frac{ 1 }{ 27 }\]A.\[\sum_{i=1}^{4}(\frac{ 1 }{ 9 })^{i-1}\]B.\[\sum_{i=1}^{4}\frac{ 3 }{ i }\]C.\[\sum_{i=1}^{4}(\frac{ 1 }{ 3 })^{i+1}\]D.\[\sum_{i=1}^{4}(\frac{ 1 }{ 3 })^{i-1}\]

OpenStudy (anonymous):

any guesses?

OpenStudy (anonymous):

looks like the denominator is powers of 3, which really only leaves the last two

OpenStudy (anonymous):

I tried plugging the numbers in and seeing if any of them matched the series but I must be doing something wrong because I can't figure it out

OpenStudy (anonymous):

okay lets try the last two, ignore the first two

OpenStudy (anonymous):

\[\sum_{i=1}^{4}(\frac{ 1 }{ 3 })^{i+1}\] put \(i=1\) and get \((\frac{1}{3})^{1+1}=(\frac{1}{3})^2=\frac{1}{9}\) as the first term

OpenStudy (anonymous):

you see right away it is wrong, because the first term should be \(1\) not \(\frac{1}{9}\)

OpenStudy (anonymous):

Yes, I follow you, so would it be D?

OpenStudy (anonymous):

now we can try \[\sum_{i=1}^{4}(\frac{ 1 }{ 3 })^{i-1}\]

OpenStudy (anonymous):

put \(i=1\) and get \((\frac{1}{3})^{1-1}=(\frac{1}{3})^0=1\) so this looks more promising

OpenStudy (anonymous):

yes, it is D you can check it works for \(i=2, 3, 4\) also

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

Thank you very much! Yes, that's it. (:

OpenStudy (anonymous):

yw